OpenStudy (anonymous):

Write equation for point slope, slope int, and standard (-5,3) parallel to y=1/2x+1/3

OpenStudy (anonymous):

m = 1/2 y -3 = 1/2 ( x +5) y = x/2 + 5/2 +3 = x/2 + 11/2 2y = x +11 -> x -2y = -11

OpenStudy (anonymous):

sorry about the signs; thanks for fixing Chlorophyll :)

OpenStudy (anonymous):

I'm glad, actually I'm double checking now!

OpenStudy (anonymous):

It takes me forever to work straight without paper and pen, also my internet connection/ PC so stuttering :(